Calculate the heat capacity of the sample beake fluid if a temperature rise from 15° to 100° C requires the absorption of 911J of heat.


1
Expert's answer
2021-06-28T08:10:50-0400


C=Q/ΔT

where,C = heat capacity,

Q = heat,

∆T = 373 -288

= 85

= 911/85

= 10.72 J/K